# TaeBench: Improving Quality of Toxic Adversarial Examples

### v1

@@ 6 Connecting to Related Works
Performing quality control of data sets used by deep learning (whether in training or during testing) is essential to ensure and enhance the overall performance and reliability of deep learning systems (Fujii et al., 2020; Wu et al., 2021; Grosman et al., 2020). (Morris et al., 2020b) proposed a set of language constraints to filter out undesirable text adversarial examples, including limits on the ratio of words to perturb, minimum angular similarity and the Part-of-Speech match constraint. The study investigated how these constraints were used to ensure the perturbation generated examples preserve the semantics and fluency of original seed text in two synonym substitution attacks against NLP classifiers. This study found the perturbations from these two attacks often do not preserve semantics, and 38% generated examples introduce grammatical errors. Two related studies from (Dyrmishi et al., 2023; Chiang and Lee, 2022) also revealed that word substitution based attack methods generate a large fraction of invalid substitution words that are ungrammatical. Both papers focus on only word substitution-based attacks attacking the general NLP classification cases, and both did not show the benefit of filtered examples.
@@
Adversarial attacks create adversarial examples designed to cause a deep learning model to make a mistake. First proposed in the image domain by (Goodfellow et al., 2014), adversarial examples provide effective lenses to measure a deep learning system’s robustness. Recent techniques that create adversarial text examples make small modifications to input text to investigate the adversarial robustness of NLP models. A body of adversarial attacks were proposed in the literature to fool question answering (Jia and Liang, 2017), machine translation (Cheng et al., 2018), text classification and more (Ebrahimi et al., 2017; Jia and Liang, 2017; Alzantot et al., 2018; Jin et al., 2019; Ren et al., 2019; Zang et al., 2020; Garg and Ramakrishnan, 2020).
@@ 7 Conclusion
In this paper, we present a model-based pipeline for quality control in in the generation of TAE. By evaluating 20+ TAE attack recipes, we curate a high-quality benchmark TaeBench. We demonstrate TaeBench’s effectiveness in assessing the robustness of real-world toxicity content moderation models, and show that adversarial training using TaeBench improves toxicity detectors’ resilience against unseen attacks.
Appendix A Limitations
Appendix B Risks and Ethical Considerations
Appendix C Appendix on Methods
C.1 Human Annotators
@@
C.2 Human Annotation of Training Data
@@ C.3 Training Configuration
C.4 On Three Local Proxy Models for Text Toxicity Detection
@@ Appendix D Appendix on Results
Table A2: Total attack time (in hours) to run 20 ToxicTrap [3] recipes. We first train a proxy target model on the train splits of each dataset, and then run ToxicTrap attacks using seeds from the test splits. Each recipe is executed using 8 Intel Xeon 2.3GHz CPUs and 1 Nvidia Tesla V100 16Gb GPU.

### v2

@@ 6 Connecting to Related Works
Performing quality control of data sets used by deep learning (whether in training or during testing) is essential to ensure and enhance the overall performance and reliability of deep learning systems (Fujii et al., 2020; Wu et al., 2021; Grosman et al., 2020). Morris et al. (2020b) proposed a set of language constraints to filter out undesirable text adversarial examples, including limits on the ratio of words to perturb, minimum angular similarity and the Part-of-Speech match constraint. The study investigated how these constraints were used to ensure the perturbation generated examples preserve the semantics and fluency of original seed text in two synonym substitution attacks against NLP classifiers. This study found the perturbations from these two attacks often do not preserve semantics, and 38% generated examples introduce grammatical errors. Two related studies from Dyrmishi et al. (2023); Chiang and Lee (2022) also revealed that word substitution based attack methods generate a large fraction of invalid substitution words that are ungrammatical. Both papers focus on only word substitution-based attacks attacking the general NLP classification cases, and both did not show the benefit of filtered examples.
@@
Adversarial attacks create adversarial examples designed to cause a deep learning model to make a mistake. First proposed in the image domain by Goodfellow et al. (2014), adversarial examples provide effective lenses to measure a deep learning system’s robustness. Recent techniques that create adversarial text examples make small modifications to input text to investigate the adversarial robustness of NLP models. A body of adversarial attacks were proposed in the literature to fool question answering (Jia and Liang, 2017), machine translation (Cheng et al., 2018), text classification and more (Ebrahimi et al., 2017; Jia and Liang, 2017; Alzantot et al., 2018; Jin et al., 2019; Ren et al., 2019; Zang et al., 2020; Garg and Ramakrishnan, 2020).
@@ 7 Conclusion
In this paper, we present a model-based pipeline for quality control in the generation of TAE. By evaluating 20+ TAE attack recipes, we curate a high-quality benchmark TaeBench. We demonstrate its effectiveness in assessing the robustness of real-world toxicity content moderation models, and show that adversarial training using TaeBench improves toxicity detectors’ resilience against unseen attacks.
Appendix A Appendix on Methods
A.1 Human Annotators
@@
A.2 Human Annotation of Training Data of TDA-BERT
@@ A.3 Training Configuration
A.4 On Three Local Proxy Models for Text Toxicity Detection
@@
Appendix B Limitations
Appendix C Risks and Ethical Considerations
@@ Appendix D Appendix on Results
Table A2: Total attack time (in hours) to run 20 ToxicTrap Bespalov et al. (2023) recipes. We first train a proxy target model on the train splits of each dataset, and then run ToxicTrap attacks using seeds from the test splits. Each recipe is executed using 8 Intel Xeon 2.3GHz CPUs and 1 Nvidia Tesla V100 16Gb GPU.